If Each of two congruent sides of an isosceles triangle is 8 in. less taht twice the base. The perimeter if the traingle is 74in. what is the length of the base?
Answer by rfer(16322)

x+2(2x-8)=74
5x-16=74
5x=90
x=18 in base
2x-8=28 in sides